A stock has produced returns of 4 percent, 10 percent, -5 percent, and 2 percent over the past four years,
respectively. What is the 99 percent probability range of returns?

-16.9 to 23.6 percent

-15.8 to 21.3 percent

-12.6 to 18.4 percent

-10.7 to 15.9 percent
Answer #1

Average return=(4%+10%-5%+2%)/4=2.75%

standard deviation using excel formula=STDEV(4%,10%,-5%,2%)=6.18%

low end of 99 percent probability range of returns=2.75%-3*6.18%=-15.8%

high end of 99 percent probability range of returns=2.75%+3*6.18%=21.3%

answer: -15.8 to 21.3 percent

the above is answer..